Write the equation of a line parallel to the given line but passing through the given point.
y=-1/3x-4 with ordered pair (-6,-3) using point slope form with work
This is as far as I have gotten I want to make sure work is correct
First substitute the slope and ordered pair
y-(-3)=1/3 (x-(-6)
Simplify the equation by eliminating the double negative and by distributing the slope
y+3=1-1/3x-(-1/3(6)
(1/3)6 =-2 we will subtract -2 from both sides
The equation for the parallel line becomes
y=-1/3x-5
The y-intercept is -5 units below the origin with the x-intercept 15 to the left of the origin.

Answer by macston(5194)

You have to solve with point slope form, but you can check your work by solving with slope-intercept form. You should have the same result.
Parallel lines have the same slope (m=-1/3), x=-6, y=-3
plug in to form:
y=mx+b
-3=-1/3(-6)+b
-3=2+b
-5=b
So the equation is:y=-1/3x-5
It checks, your done.